# Random Lévy Looptrees and Lévy Maps

### v2

@@ 2.2.   Lévy looptrees and Gaussian labels
Recall from Section 1.2 the model of Lévy processes that we consider: $X$ has no negative jump, infinite variation paths, and its characteristic function is integrable (1.3). Let us mention that, this integrability condition holds as soon as $\beta>0$ and also when $u^{-2}|\log u|^{-1}\int_{0}^{u}r^{2}\pi(\mathop{}\!\mathrm{d}r)\to\infty$ as $u\to 0$, where $\pi$ is the Lévy measure of $X$, see [Kni96, Section 2], which is taken from [Kal81, Section 5]. By inverse Fourier transform, the process $X$ then admits transition densities, say $(p_{t}(x))_{t>0,x\in\mathbb{R}}$, which are jointly continuous in time and space as shown e.g. in [Kni96, Lemma 2.4]. In addition, they never vanish, and precisely Sharpe [Sha69] shows that if one of them vanishes somewhere, then the process $X$ is, up to a drift, either a subordinator or the negative of one.
@@
Let $Y$ be any càdlàg function on $[0,\infty)$ with no negative jump, which can be e.g. the Lévy process $X$ or its excursion $X^{\mathrm{ex}}$ extended to 00 on $[1,\infty)$. Let us recall from [Mar22b] the construction of the looptree coded by $Y$ as well as random Gaussian labels on it. This originates from [LGM11, CK14] which applies in a pure jump case, where the continuous part $C$ below vanishes. First define from $Y$ a partial order $\preceq$ by setting for every $s,t\geq 0$:
@@
where ||$\cdot$|| ⋅ | stands for the Lebesgue measure. It is standard that the function defined for every $0\leq s\leq t\leq 1$ by:
$d_{C}(s,t)=d_{C}(t,s)=C_{s}+C_{t-}-2\inf_{[s,t]}C$ (2.5)
is a pseudo-distance and the quotient space $T_{C}$ obtained by identifying all pairs of times at $d_{C}$-distance 00 is a compact rooted real tree [LG05].
@@
We may now define the looptree coded by $Y$. We associate with any jump time $t\geq 0$ a cycle $[0,\Delta Y_{t}]$ with $\Delta Y_{t}$ identified with 00, equipped with the metric $\delta_{t}(a,b)=\min\{|a-b|,\Delta Y_{t}-|a-b|\}$ for all $a,b\in[0,\Delta Y_{t}]$. For definiteness, if $\Delta Y_{t}=0$, then we set $\delta_{t}(0,0)=0$. Define the looptree distance with parameter $a\geq 0$ between any $s,t\geq 0$ by:
@@
It can be checked that $d^{a}_{\mathscr{L}(Y)}$ is indeed a continuous pseudo-distance on $[0,1]$ for every $a$, which moreover satisfies:
@@
where the right-hand side is defined in (2.5), see [Mar22b, Proposition 3.2] which extends [CK14, Lemma 2.1] when $C=0$. Henceforth we denote by $\mathscr{L}^{a}(Y)$ the quotient space obtained by identifying all pairs of times at pseudo-distance 00. We shall simply write $\mathscr{L}(Y)$ and $d_{\mathscr{L}(Y)}$ for $\mathscr{L}^{1}(Y)$ and $d^{1}_{\mathscr{L}(Y)}$. The value of $a>0$ does not affect the fractal dimensions of the looptree, and Theorem 1.1 extends to $\mathscr{L}^{a}(X^{\mathrm{ex}})$. This constant only appears in the invariance principles in Section 2.3. In the case where $Y$ is a Brownian excursion, we have $C=Y$ so $\mathscr{L}^{a}(Y)=a\cdot T_{C}=a\cdot T_{Y}$ is a scaled Brownian tree coded by $Y$.
@@
We construct a Gaussian field on the looptree $\mathscr{L}^{a}(Y)$ by placing on each cycle an independent Brownian bridge, with duration given by the length of the cycle, which describes the increments of the field along the cycle. Formally, recall that the standard Brownian bridge $b=(b_{t})_{t\in[0,1]}$ is a centred Gaussian process with covariance
$\mathop{}\!\mathbb{E}[b_{s}b_{t}]=\min(s,t)-st,\qquad s,t\in[0,1].$
One can consider a bridge of any given duration using a diffusive scaling. Now recall the partial order $\prec$ as well as the notation $R^{t}_{s}=\inf_{r\in[s,t]}Y_{r}-Y_{s-}$ from (2.3). Let $C$ be the continuous part defined in (2.4) and let $Z^{C}$ be associated Brownian snake. Independently let $(b_{i})_{i\geq 1}$ denote a sequence of i.i.d. standard Brownian bridges, and define for every $a\geq 0$ and $t\in[0,1]$,
@@
$\mathop{}\!\mathbb{E}\biggl{[}\Bigl{(}\sqrt{\Delta Y_{t_{i}}}b_{i}\bigl{(}(\Delta Y_{t_{i}})^{-1}R^{t}_{t_{i}}\bigr{)}\Bigr{)}^{2}\biggr{]}=\frac{R^{t}_{t_{i}}(\Delta Y_{t_{i}}-R^{t}_{t_{i}})}{\Delta Y_{t_{i}}}\leq\min\{R^{t}_{t_{i}},\Delta Y_{t_{i}}-R^{t}_{t_{i}}\}=\delta_{t_{i}}(0,R^{t}_{t_{i}}).$
@@
Taking $s=0$, this shows that the random variable $Z^{a}_{t}$ in (2.8) is well-defined for any fixed $t$, which in particular implies that the series in (2.8) converges. Further, given $Y$, we see that any pair $s,t\geq 0$ that has $d^{a}_{\mathscr{L}(Y)}(s,t)=0$ also has $Z^{a}_{s}=Z^{a}_{t}$ almost surely so $Z^{a}$ can actually be seen as a random process indexed by the looptree. Note that by the scaling property of Gaussian variables the process $cZ^{a}$ constructed from $Y$ has the same law as the process $Z^{a}$ but constructed from $c^{2}Y$. The value $a=1/3$ shall play a predominant role in relation with maps and we simply write $Z=Z^{1/3}$. We shall also denote by $Z^{\mathrm{ex}}$ this process $Z=Z^{1/3}$ constructed from
@@ 2.3.   Invariance principles
We have not defined Lévy maps yet; they will shortly be defined as subsequential limits of rescaled finite maps. Before that, let us start by discussing limits of labelled looptrees built from random bridges with exchangeable increments. Typical examples of applications include the Boltzmann distributions recalled earlier, which relate to random walk excursions of duration $n$, with a step distribution which may depend on $n$. We let $X$ be a Lévy process with Laplace exponent $\psi$, with Gaussian parameter $\beta\geq 0$, and we let $X^{\mathrm{br}}$ and $X^{\mathrm{ex}}$ denote the unit duration bridge and excursion versions of $X$ respectively. For $a\geq 0$, we also let $\mathscr{L}^{a}(X^{\mathrm{ex}})$ denote the looptree defined as in (2.6) and the lines below. Finally, let $Z^{\mathrm{ex}}$ denote the label process $Z^{a}$ defined as in (2.8) with $a=1/3$ and $Y=X^{\mathrm{ex}}$. The next result is an easy consequence of [Mar22b, Theorems 7.4 and 7.9] by conditioning on the increment sizes.
@@
For every $n\geq 1$, let $B^{n}=(B^{n}_{k})_{0\leq k\leq n+1}$ be a random bridge, with $B^{n}_{n+1}=-1$, with increments $\Delta B^{n}_{i}\in\mathbb{Z}_{\geq-1}$ and which are exchangeable. Let $W^{n}=(W^{n}_{k})_{0\leq k\leq n+1}$ denote the excursion obtained by cyclicly shifted $B^{n}$ at its first minimum. Let further $Z^{n}$ denote the label process associated with a uniform random good labelling of the looptree coded by $W^{n}$.
@@
    where $a=0$ by convention when $\beta=0$. Then we have the convergence in distribution:
    $r_{n}^{-1}\,\mathsf{L}(W^{n})\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip\mathscr{L}^{(a+1)/2}(X^{\mathrm{ex}})$
    for the Gromov–Hausdorff–Prokhorov topology where $\mathsf{L}(W^{n})$ is equipped either with the uniform probability on the vertices or the uniform probability on the corners.
@@
Let us comment on Assumption (2.10). When following a geodesic path in the looptree and traversing a cycle, one takes the shortest between the left and right length. Roughly speaking, typically half of the cycles are traversed on the right and half on the left, thus the microscopic cycles contributes to roughly half the continuous part of the coding path, leading to $\mathscr{L}^{1/2}(X^{\mathrm{ex}})$ in the limit. However, at the discrete level, the fact that the cycle has even or odd length influences the law of the length of the smallest part between left and right, namely $\sum_{j=1}^{k}\min(j,k-j)$ depends on the parity of $k$. This contribution is precisely captured in (2.10) and adds an extra factor to the continuous part in the looptree distance in the limit. Such a phenomenon was already observed in [KR20, Theorem 1.2] (see however [Mar22b, Remark 7.7] on the extra constant). We note that this on the other hand has no influence on the labels. The parameter $1/3$ appears in the limit due to our choice of labels in the discrete model, see [Mar22b, Theorem 7.9] for more general labels.
@@
$\theta_{i}=\Bigl{(}\beta^{2}+\sum_{i\geq 1}(\Delta X^{\shortdownarrow}_{i})^{2}\Bigr{)}^{-1/2}\Delta X^{\shortdownarrow}_{i}\qquad\text{and then}\qquad\theta_{0}^{2}=1-\sum_{i\geq 1}\theta_{i}^{2}=\frac{\beta^{2}}{\beta^{2}+\sum_{i\geq 1}(\Delta X^{\shortdownarrow}_{i})^{2}}.$
@@
which is the exchangeable increment process with Gaussian parameter $\theta_{0}$ and jump sizes $(\theta_{i})_{i\geq 1}$. Then the assumptions of [Mar22b, Theorem 7.4] are fulfilled, namely let $\sigma^{2}_{n}=\sum_{i=1}^{n}\Delta B^{n}_{i}(\Delta B^{n}_{i}+1)=\sum_{i=1}^{n}(\Delta B^{n}_{i})^{2}-1$, then according to [Kal02, Theorem 16.23] the convergence of $r_{n}^{-1}B^{n}$ to $X^{\mathrm{br}}$ is equivalent to the convergence for every $i\geq 1$ of the $i$’th largest increment of $r_{n}^{-1}B^{n}$ to $\Delta X^{\shortdownarrow}_{i}$ and the convergence $r_{n}^{-2}\sigma_{n}^{2}\to\beta^{2}+\sum_{i\geq 1}(\Delta X^{\shortdownarrow}_{i})^{2}$. Thus if we instead rescale $B^{n}$ by $\sigma_{n}^{-1}$, then for every $i\geq 1$, the $i$’th largest increment of $\sigma_{n}^{-1}B^{n}$ converges to $(\beta^{2}+\sum_{i\geq 1}(\Delta X^{\shortdownarrow}_{i})^{2})^{-1/2}\Delta X^{\shortdownarrow}_{i}=\theta_{i}$, while
@@
We can then apply [Mar22b, Theorem 7.4] and deduce that.
$r_{n}^{-1}\,\mathsf{L}(W^{n})\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip\Bigl{(}\beta^{2}+\sum_{i\geq 1}(\Delta X^{\shortdownarrow}_{i})^{2}\Bigr{)}^{1/2}\mathscr{L}^{(a+1)/2}(\mathcal{V}Y)=\mathscr{L}^{(a+1)/2}(X^{\mathrm{ex}}),$
where the equality simply follows from the identity $c\cdot\mathscr{L}^{(a+1)/2}(\mathcal{V}Y)=\mathscr{L}^{(a+1)/2}(c\mathcal{V}Y)$ for every $c>0$, which is plain from the construction of the looptree distance (2.6).
As for the labels, we apply [Mar22b, Theorem 7.9] in the particular case of the labels that code maps (see the beginning of [Mar22b, Section 7.4]) and deduce the convergence in distribution
@@
Remark 2.3.
Let us mention that behind the abstract Gromov–Hausdorff–Prokhorov convergence of the looptrees lies a very explicit convergence. Indeed, let us denote by $(v^{n}_{k})_{0\leq k\leq n}$ the vertices of $\mathsf{L}(W^{n})$ encountered when following its contour. Formally, recall from Section 2.1 the sequence $(e_{0},\dots,e_{n})$ of oriented edges (with $e_{n}=e_{0}$), then set $v^{n}_{k}$ to be the origin of $e_{k}$. One can then define a pseudo-distance $d^{n}$ on $\{0,\dots,n\}$ by letting $d^{n}(i,j)$ be the looptree distance between $v^{n}_{i}$ and $v^{n}_{j}$. It is shown in [Mar22b] that
$(r_{n}^{-1}d^{n}(\lfloor ns\rfloor,\lfloor nt\rfloor))_{s,t\in[0,1]}\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip(d^{(a+1)/2}_{\mathscr{L}(X^{\mathrm{ex}})}(s,t))_{s,t\in[0,1]}$
for the uniform topology. This immediately implies the Gromov–Hausdorff–Prokhorov convergence for the uniform law on the corners. In order to control the uniform law on the vertices, one needs to get rid of the redundancies in the list $(v^{n}_{0},\dots,v^{n}_{n})$, and the argument is provided by [Mar22a, Lemma 2.2] with $A=\{-1\}$.
Theorem 2.1 finds applications to scaling limits of random planar maps as in [Mar22b, Theorem 7.12]. Indeed, recall that a pointed bipartite map $(M^{n},v_{\star}^{n})$ with $n$ edges is coded by a labelled looptree such that the labels, shifted to have minimum $1$, code the graph distances to $v^{n}_{\star}$ in $M^{n}$. Let us denote by $(v^{n}_{k})_{0\leq k\leq n}$ the vertices of the looptree encountered when following its contour as in the previous remark. They can naturally be seen through the bijection as the vertices of the map $M^{n}$ different from $v_{\star}^{n}$; in particular $v^{n}_{0}$ is an extremity of the root edge of $M^{n}$ (the one farther away from $v_{\star}^{n}$). Finally let $d_{M^{n}}$ denote the graph distance in the map $M^{n}$.
@@
Let $(M^{n},v_{\star}^{n})$ be a pointed map with $n$ edges and assume that the Łukasiewicz path $W^{n}$ of its associated looptree satisfies the assumptions of Theorem 2.1. Then the following holds jointly with this theorem.
@@
    $\Bigl{(}\frac{1}{\sqrt{2r_{n}}}d_{M^{n}}(v^{n}_{\star},v^{n}_{\lfloor nt\rfloor})\Bigr{)}_{t\in[0,1]}\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip(Z^{\mathrm{ex}}_{t}-\min Z^{\mathrm{ex}})_{t\in[0,1]},$
    holds for the uniform topology. Consequently, if $V(M^{n})$ denotes the set of vertices of $M^{n}$, then
    $\frac{1}{\sqrt{2r_{n}}}d_{M^{n}}(v^{n}_{\star},v^{n}_{0})\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip-\min Z^{\mathrm{ex}}\quad\text{and}\quad\frac{1}{\sqrt{2r_{n}}}\max_{v\in V(M^{n})}d_{M^{n}}(v^{n}_{\star},v)\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip\max Z^{\mathrm{ex}}-\min Z^{\mathrm{ex}},$
    and for every continuous and bounded function $F$, we have:
    $\frac{1}{\#V(M^{n})}\sum_{v\in V(M^{n})}F\Bigl{(}\frac{1}{\sqrt{2r_{n}}}d_{M^{n}}(v^{n}_{\star},v)\Bigr{)}\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip\int_{0}^{1}F(Z^{\mathrm{ex}}_{t}-\min Z^{\mathrm{ex}})\mathop{}\!\mathrm{d}t.$

### v3

@@ 2.2.   Lévy looptrees and Gaussian labels
Recall from Section 1.2 the model of Lévy processes that we consider: $X$ has no negative jump, infinite variation paths, and its characteristic function is integrable (1.3). By inverse Fourier transform, the process $X$ then admits transition densities, say $(p_{t}(x))_{t>0,x\in\mathbb{R}}$, which are jointly continuous in time and space as shown e.g. in [Kni96, Lemma 2.4]. In addition, they never vanish, and precisely Sharpe [Sha69] shows that if one of them vanishes somewhere, then the process $X$ is, up to a drift, either a subordinator or the negative of one.
@@
Let $Y$ be any càdlàg function on $[0,\infty)$ with no negative jump, which can be e.g. the Lévy process $X$ or its excursion $X^{\mathrm{ex}}$ extended to 00 on $[1,\infty)$. Let us recall from [Mar24] the construction of the looptree coded by $Y$ as well as random Gaussian labels on it. The reader may alternatively find this construction of looptrees in [Kha22] where they are called ‘vernation trees’, together with additional topological details on these objects. This originates from [LGM11, CK14] which applies in a pure jump case, where the continuous part $C$ below vanishes. First define from $Y$ a partial order $\preceq$ by setting for every $s,t\geq 0$:
@@
where ||$\cdot$|| ⋅ | stands for the Lebesgue measure. For a càdlàg function $g$ on $[0,\infty)$, with no negative jump, let us define for every $0\leq s\leq t$:
$d_{g}(s,t)=d_{g}(t,s)=g_{s}+g_{t-}-2\inf_{[s,t]}g.$ (2.5)
Note that we take the left-limit at time $t$. When $g$ is continuous, it is well-known that $d_{g}$ is a pseudo-distance and that the quotient space $T_{g}$ obtained by identifying all pairs of times at $d_{g}$-distance 00 is a rooted real tree [LG05]. In particular one may define a tree $T_{C}$ from the continuous part $C$ from (2.4). We shall also use $d_{g}$ for $g=Y$ to upper bound the looptree distance, see Equation (2.7) below.
@@
We may now define the looptree coded by $Y$. We associate with any jump time $t\geq 0$ a cycle $[0,\Delta Y_{t}]$ with $\Delta Y_{t}$ identified with 00, equipped with the metric $\delta_{t}(a,b)=\min\{|a-b|,\Delta Y_{t}-|a-b|\}$ for all $a,b\in[0,\Delta Y_{t}]$. For definiteness, if $\Delta Y_{t}=0$, then we set $\delta_{t}(0,0)=0$. Following [Mar24], define the looptree distance with parameter $a\geq 0$ between any $s,t\geq 0$ by:
@@
It can be checked that $d^{a}_{\mathscr{L}(Y)}$ is indeed a continuous pseudo-distance on $[0,1]$ for every $a$, which moreover satisfies the following inequality: let $d_{Y}$ be defined as $d_{g}$ in (2.5) with $g=Y$, then
@@
see [Mar24, Proposition 3.2] which extends [CK14, Lemma 2.1] when $C=0$. Henceforth we denote by $\mathscr{L}^{a}(Y)$ the quotient space obtained by identifying all pairs of times at pseudo-distance 00. We shall simply write $\mathscr{L}(Y)$ and $d_{\mathscr{L}(Y)}$ for $\mathscr{L}^{1}(Y)$ and $d^{1}_{\mathscr{L}(Y)}$. The value of $a>0$ does not affect the fractal dimensions of the looptree, and Theorem 1.1 extends to $\mathscr{L}^{a}(X^{\mathrm{ex}})$. This constant only appears in the invariance principles in Section 2.3. In the case where $Y$ is a Brownian excursion, we have $C=Y$ so $\mathscr{L}^{a}(Y)=a\cdot T_{C}=a\cdot T_{Y}$ is a scaled Brownian tree coded by $Y$. The metric space $\mathscr{L}^{a}(Y)$ has been introduced in [Mar24], and in the particular case $a=1/2$ in [Kha22] it is called the vernation tree coded by $Y$.
@@
We construct a Gaussian field on the looptree $\mathscr{L}^{a}(Y)$ by placing on each cycle an independent Brownian bridge, with duration given by the length of the cycle, which describes the increments of the field along the cycle. Formally, recall that the standard Brownian bridge $b=(b(t))_{t\in[0,1]}$ is a centred Gaussian process with covariance
$\mathop{}\!\mathbb{E}[b(s)b(t)]=\min(s,t)-st,\qquad s,t\in[0,1].$
One can consider a bridge of any given duration using a diffusive scaling. Now recall the partial order $\prec$ as well as the notation $R^{t}_{s}=\inf_{r\in[s,t]}Y_{r}-Y_{s-}$ from (2.3). Let $C$ be the continuous part defined in (2.4) and let $Z^{C}$ be the associated Brownian snake. Independently let $(b_{i})_{i\geq 1}$ denote a sequence of i.i.d. standard Brownian bridges and define for every $a\geq 0$ and $t\in[0,1]$:
@@
$\mathop{}\!\mathbb{E}\biggl{[}\Bigl{(}\sqrt{\Delta Y_{t_{i}}}b_{i}\bigl{(}(\Delta Y_{t_{i}})^{-1}R^{t}_{t_{i}}\bigr{)}\Bigr{)}^{2}\biggr{]}$ $=\frac{R^{t}_{t_{i}}(\Delta Y_{t_{i}}-R^{t}_{t_{i}})}{\Delta Y_{t_{i}}}$
$\leq\min\{R^{t}_{t_{i}},\Delta Y_{t_{i}}-R^{t}_{t_{i}}\}=\delta_{t_{i}}(0,R^{t}_{t_{i}}).$
@@
Taking $s=0$, this shows that the random variable $Z^{a}_{t}$ in (2.8) is well-defined for any fixed $t$, which in particular implies that the series in (2.8) converges. Further, given $Y$, we see that any pair $s,t\geq 0$ that has $d^{a}_{\mathscr{L}(Y)}(s,t)=0$ also has $Z^{a}_{s}=Z^{a}_{t}$ almost surely so $Z^{a}$ can actually be seen as a random process indexed by the looptree. Note that by the scaling property of Gaussian variables, the process $cZ^{a}$ constructed from $Y$ has the same law as the process $Z^{a}$ but constructed from $c^{2}Y$. The value $a=1/3$ shall play a predominant role in relation with maps and we simply write $Z=Z^{1/3}$. We shall also denote by $Z^{\mathrm{ex}}$ this process $Z=Z^{1/3}$ constructed from $Y=X^{\mathrm{ex}}$.
@@ 2.3.   Invariance principles
We have not defined Lévy maps yet; they will shortly be defined as subsequential limits of rescaled finite random maps. Precisely, we shall consider rooted bipartite maps $M^{n}$ with $n$ edges that carry a distinguished vertex $v^{n}_{0}$ and we shall assume that their distribution for every $n$ only depends on their face degrees in the sense that if $(m,v)$ and $(m^{\prime},v^{\prime})$ are two (rooted bipartite and) pointed maps with $n$ edges that have the same amount of faces with degree $2k$ for every $k\geq 1$, then
$\mathop{}\!\mathbb{P}((M^{n},v^{n}_{0})=(m,v))=\mathop{}\!\mathbb{P}((M^{n},v^{n}_{0})=(m^{\prime},v^{\prime})).$ (2.10)
Typical examples include the conditioned Boltzmann distributions recalled earlier, since $(m,v)$ and $(m^{\prime},v^{\prime})$ have the same weight. We may here generalise this distribution by letting the weight sequence $q$ of the Boltzmann law vary with $n$.
This assumption has two consequences. First, conditionally on its number of faces of each degree $(M^{n},v^{n}_{0})$ has the uniform distribution on maps with these degree statistics, which will falls in the scope of [Mar22, Mar24]. Second, conditionally on $M^{n}$, the distinguished vertex $v^{n}_{0}$ has the uniform distribution on the vertices of $M^{n}$. This will only be used in this work in order to identify the limit as the Brownian sphere or a stable map in the appropriate regime.
By construction of the bijection presented earlier between pointed maps and well-labelled looptrees, and further with pairs of processes, if the random map $(M^{n},v^{n}_{0})$ satisfies (2.10), then its associated pair $(W^{n},Z^{n})$ gives the same probability to all pairs in which the Łukasiewicz paths have the same amount of jumps of size $k$ for every $k\geq-1$. In particular, conditionally on its increment sizes, the path $W^{n}$ is an excursion sampled uniformly at random among all possibilities, and $Z^{n}$ is further obtained conditionally given $W^{n}$ as a uniformly chosen random good labelling of the looptree, obtained by placing independent uniform random bridges around each cycles. By the so-called cycle lemma, the law of $W^{n}$ is that of the path obtained from a random bridge $B^{n}=(B^{n}_{k})_{0\leq k\leq n+1}$, with $B^{n}_{n+1}=-1$, and with increments $\Delta B^{n}_{i}\in\mathbb{Z}_{\geq-1}$ that are exchangeable, by cyclicly shifting $B^{n}$ at its first minimum to turn it into an excursion.
Let us next say a few words on the notion of convergence that we consider. The abstract framework to study scaling limit of (possibly pointed) finite graphs, such as (loop)trees and maps that has been developed is that of the Gromov–Hausdorff–Prokhorov (GHP) distance. The latter is a separable and complete distance on the space of isometry classes of compact measured metric spaces, possibly with distinguished points. Then we say that rescaled graphs converge when their set of vertices equipped with the rescaled graph distance and the uniform probability measure do in the GHP sense. In our case, we can make this convergence much more explicit. Indeed, recall from Section 2.1 that the vertices of looptrees and maps can ordered in some canonical way $(v^{n}_{0},\dots,v^{n}_{N})$ where $n$ is the number of edges and $N+1$ the number of vertices of the map, and where for simplicity we let $v^{n}_{0}=v^{n}_{N}$ in the case of the looptree, whereas in the map $v^{n}_{0}$ is the distinguished vertex. One can then encode the graph distances in each model into a function $d^{n}$ on $\{0,\dots,N\}$ by letting $d^{n}(i,j)$ be the distance between $v^{n}_{i}$ and $v^{n}_{j}$ in the corresponding model, which we transform into a function on $[0,1]^{2}$ by a bilinear interpolation and scaling by $N$. Then if for some scaling factor $r_{n}\to\infty$ we have
$(r_{n}^{-1}d^{n}(s,t))_{s,t\in[0,1]}\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip(d(s,t))_{s,t\in[0,1]}$
for the uniform topology, for some continuous pseudo-distance $d$, then the graph rescaled by the factor $r_{n}^{-1}$ indeed converges for the GHP distance to the quotient space $[0,1]/\{d=0\}$ obtained by identifying all pairs of times at distance 00. This convergence is actually what is proved for Brownian and stable looptrees and maps [LG13, Mie13, CK15, CMR25].
@@
Let us apply this to random looptrees first and then to maps. We let $X$ be a Lévy process with Laplace exponent $\psi$, with Gaussian parameter $\beta\geq 0$, and we let $X^{\mathrm{br}}$ and $X^{\mathrm{ex}}$ denote the unit duration bridge and excursion versions of $X$ respectively. For $a\geq 0$, we also let $\mathscr{L}^{a}(X^{\mathrm{ex}})$ denote the looptree defined as in (2.6) and the lines below. Finally, let $Z^{\mathrm{ex}}$ denote the label process $Z^{a}$ defined as in (2.8) with $a=1/3$ and $Y=X^{\mathrm{ex}}$. The next result is an easy consequence of [Mar24, Theorems 7.4 and 7.9] by conditioning on the increment sizes. Below $d_{\mathsf{L}(W^{n})}$ denotes the function on $[0,1]^{2}$ that encodes the graph distances in the looptree $\mathsf{L}(W^{n})$ as introduced above.
@@
For every $n\geq 1$, let $B^{n}=(B^{n}_{k})_{0\leq k\leq n+1}$ be a random bridge, with $B^{n}_{n+1}=-1$, with increments $\Delta B^{n}_{i}\in\mathbb{Z}_{\geq-1}$ and which are exchangeable. Let $W^{n}=(W^{n}_{k})_{0\leq k\leq n+1}$ denote the excursion obtained by cyclicly shifting $B^{n}$ at its first minimum. Let further $Z^{n}$ denote the label process associated with a uniform random good labelling of the looptree coded by $W^{n}$.
@@
    where $a=0$ by convention when $\beta=0$. Then, jointly with the first part, we have:
    $r_{n}^{-1}\,d_{\mathsf{L}(W^{n})}\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip d^{(a+1)/2}_{\mathscr{L}(X^{\mathrm{ex}})},$
    in the space of real-valued continuous functions defined on $[0,1]^{2}$ equipped with the uniform topology. This implies that $r_{n}^{-1}\,\mathsf{L}(W^{n})$ converges in distribution to $\mathscr{L}^{(a+1)/2}(X^{\mathrm{ex}})$ for the Gromov–Hausdorff–Prokhorov topology.
Let us mention that [Kha22, Section 6.3] obtains limit theorems for looptrees coded by excursions, but there are several major differences. First, as we have already mentioned in the beginning of Sec. 2.1, the discrete looptree coded $W^{n}$ considered here slightly differs from the version studied in [Kha22]. Second, in our setting $X$ is not necessarily pure jump and we may have $a\neq 1/2$, which is outside of the scope of vernation trees considered in [Kha22].
@@
Let us comment on Assumption (2.11). When following a geodesic path in the looptree and traversing a cycle, one takes the shortest between the left and right length. Roughly speaking, typically half of the cycles are traversed on the right and half on the left, thus the microscopic cycles contributes to roughly half the continuous part of the coding path, leading to $\mathscr{L}^{1/2}(X^{\mathrm{ex}})$ in the limit. However, at the discrete level, the fact that the cycle has even or odd length influences the law of the length of the smallest part between left and right, namely $\sum_{j=1}^{k}\min(j,k-j)$ depends on the parity of $k$. This contribution is precisely captured in (2.11) and adds an extra factor to the continuous part in the looptree distance in the limit. Such a phenomenon was already observed in [KR20, Theorem 1.2] (see however [Mar24, Remark 7.7] on the extra constant). We note that this has however no influence on the labels. The parameter $1/3$ appears in the limit due to our choice of labels in the discrete model, see [Mar24, Theorem 7.9] for more general labels.
@@
$\theta_{i}=\Bigl{(}\beta^{2}+\sum_{i\geq 1}(\Delta X^{\shortdownarrow}_{i})^{2}\Bigr{)}^{-1/2}\Delta X^{\shortdownarrow}_{i}\quad\text{and then}\quad\theta_{0}^{2}=1-\sum_{i\geq 1}\theta_{i}^{2}=\frac{\beta^{2}}{\beta^{2}+\sum_{i\geq 1}(\Delta X^{\shortdownarrow}_{i})^{2}}.$
@@
be the exchangeable increment process with Gaussian parameter $\theta_{0}$ and jump sizes $(\theta_{i})_{i\geq 1}$. Then the assumptions of [Mar24, Theorem 7.4] are fulfilled, namely let $\sigma^{2}_{n}=\sum_{i=1}^{n}\Delta B^{n}_{i}(\Delta B^{n}_{i}+1)=\sum_{i=1}^{n}(\Delta B^{n}_{i})^{2}-1$, then according to [Kal02, Theorem 16.23] the convergence of $r_{n}^{-1}B^{n}$ to $X^{\mathrm{br}}$ is equivalent to the convergence for every $i\geq 1$ of the $i$’th largest increment of $r_{n}^{-1}B^{n}$ to $\Delta X^{\shortdownarrow}_{i}$ and the convergence $r_{n}^{-2}\sigma_{n}^{2}\to\beta^{2}+\sum_{i\geq 1}(\Delta X^{\shortdownarrow}_{i})^{2}$. Thus if we instead rescale $B^{n}$ by $\sigma_{n}^{-1}$, then for every $i\geq 1$, the $i$’th largest increment of $\sigma_{n}^{-1}B^{n}$ converges to $(\beta^{2}+\sum_{i\geq 1}(\Delta X^{\shortdownarrow}_{i})^{2})^{-1/2}\Delta X^{\shortdownarrow}_{i}=\theta_{i}$, while
@@
It is plain from the construction of the looptree distance (2.6) that for every $c>0$, we have $c\cdot\mathscr{L}^{(a+1)/2}(\mathcal{V}Y)=\mathscr{L}^{(a+1)/2}(c\mathcal{V}Y)$. Then applying [Mar24, Theorem 7.4], or precisely Equation 38 there, we deduce that
$r_{n}^{-1}\,d_{\mathsf{L}(W^{n})}\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip\Bigl{(}\beta^{2}+\sum_{i\geq 1}(\Delta X^{\shortdownarrow}_{i})^{2}\Bigr{)}^{1/2}d^{(a+1)/2}_{\mathscr{L}(\mathcal{V}Y)}=d^{(a+1)/2}_{\mathscr{L}(X^{\mathrm{ex}})},$
for the uniform topology.
As for the labels, we apply [Mar24, Theorem 7.9] in the particular case of the labels that code maps (see the beginning of [Mar24, Section 7.4]) and deduce the convergence in distribution
@@
Theorem 2.1 finds applications to scaling limits of random planar maps which is now quite classical, and we refer to e.g. [Mar24, Section 2.4] for a recent account. Let $d_{M^{n}}$ denote the function on $[0,1]^{2}$ that encodes the graph distances in the map $M^{n}$ induced by the order on the vertices inherited from that on the looptree. Recall that in this order, the first vertex is the distinguished one $v_{0}^{n}$, while the last vertex is the extremity of the root edge that is the farthest away from $v_{0}^{n}$.
@@
Let $(M^{n},v_{0}^{n})$ be a random pointed map with $n$ edges whose distribution satisfies (2.10). Let $W^{n}$ and $Z^{n}$ denote its associated Łukasiewicz path and label process and assume that they satisfy the assumptions of Theorem 2.1. Then the following holds jointly with this theorem.
@@
    $\Bigl{(}\frac{1}{\sqrt{2r_{n}}}d_{M^{n}}(0,t)\Bigr{)}_{t\in[0,1]}\enskip\mathop{\longrightarrow}^{(d)}_{n\to\infty}\enskip(Z^{\mathrm{ex}}_{t}-\min Z^{\mathrm{ex}})_{t\in[0,1]},$
    holds for the uniform topology.